# A way around the exponential scaling in optimal quantum control

### v1

@@ IV Realization of propagators
As sketched in Sec. II.3.2, the current approach can also be used to construct a time-dependent Hamiltonian to induce dynamics that coincides with a desired unitary $U_{T}$ at a given time $T$. This is exemplified here in more detail for $U_{T}=\exp(-i\pi/8\,H_{T})$ with $H_{T}$ given by the three Hamiltonians $H_{G}$ (Eq. (22)), $H_{C}$ (Eq. (23)) and $H_{D}$ (Eq. (24)), and system dynamics in terms of the Hamiltonian $H(t)$ in Eq. (20).
@@ IV.1 Objective function
As shown in Sec. A, the only operators that commute with each of the operators in Eq. (32c) are multiples of the identity, which leaves $U_{T}^{\dagger}U=\exp(i\varphi)\mathbbm{1}$ as the only solution.
@@ IV.2 Results – Propagators
Fig. 2b depicts the infidelities ${\cal J}_{{\cal U}_{T}}$ defined in Eq. (14) together with Eq. (32c) for optimized Hamiltonian as function of system size $n$ with solid red diamonds for $U_{T}=\exp(-i\pi/8\,H_{G})$, solid blue squares for $U_{T}=\exp(-i\pi/8\,H_{C})$ and solid black triangles for $U_{T}=\exp(-i\pi/8\,H_{D})$.
@@
Also the duration of controlled dynamics required for the realization of a desired propagator, has a fundamental limit given by the interaction strength $g$. The numerically observed minimal time required to implement the unitaries induced by $H_{C}$ and $H_{D}$ is independent of the system size ($\sim\pi/2g$), but the minimal time required to implement the unitary induced by $H_{G}$ scales linearly in the number of qubits ($\sim 2n\pi/g$). Given that $H_{C}$ and $H_{D}$ contain at most three-body interactions, whereas $H_{G}$ contains an $n$-body interaction, i.e. an interaction whose type changes with increasing system size $n$, the numerically observed dependence of minimal durations of controlled dynamics seem consistent with the underlying problem, which generates some confidence that the obtained solutions are close to the actual global solution of the control problem.
@@ V.2 Driven spin comb
Figure 3: Spin comb with nearest-neighbour $XY$ interactions in the top chain and $ZZ$ interactions between spins of different chains (inset (a)). Single-spin $X$-driving (inset (a)) enables the realization of effective dynamics with four-spin interactions matching the Toric Code Plaquette operators Eq. (42) (inset (b)).
An interaction geometry that results in a Lie algebra of quadratic size in the qubit number $n$ that can be used to realize elementary building blocks of topologically protected quantum information processing [55] is depicted in Fig. 3(a). It is a comb-structure with a linear chain of $m=n/2$ qubits with nearest-neighbour interactions, and a second chain of another $m$ qubits interacting only with qubits on the first chain.
@@ V.3 Hexagonal spin ladder
with $m=n/2$, $j\in[1,(m-1)/2]$, $l\in[1,(m+1)/2]$. The resultant Lie algebra can be constructed systematically similarly to the discussion in Sec. V.1; because of the more intricate interaction geometry the exact form of the individual elements is a bit more bulky, and it is deferred to Appendix C.
@@
VI Outlook
While the exponential scaling of composite quantum systems poses fundamental constraints that can not be overcome in general, the fact that there are invariants with favourable scaling does allow us to overcome them in specific cases. With control over interaction geometries in various platforms such as superconducting qubits [59], trapped ions and Rydberg atoms [60] the present approach can help to fully exploit the qubit count in current devices (ranging up to $256$ [60]) without the need of long gate sequences that are conflicting with decoherence and the accumulation of individual gate errors. While verification of the functionality of a quantum device is a challenging problem [61], the description of quantum dynamics in terms of invariants can also be used to make predictions for the targeted dynamics that can be experimentally tested.
@@ Acknowledgements
Appendix A Absence of Symmetries in the Control Hamiltonian
@@
$H(t)=\sum_{j=1}^{n}f_{j}(t)Z_{j}+g\sum_{j=1}^{n-1}X_{j}X_{j+1}+\sum_{j\in\{1,n\}}h_{j}(t)X_{j}$ (45)
for $g\neq 0$ and general time-dependent functions $f_{j}(t)$ and $h_{j}(t)$, i.e.
@@
with an $n-2$-qubit operator $D^{n}$ that can be expanded in terms of tensor-products of $\mathbbm{1}$ and $Z$ terms, or, more explicitely as
@@
Appendix B Numerical optimisation
@@
This section describes the numerical techniques used to find the optimal control strategies for state transfer and propagator realization problems, as detailed in Secs. III and IV. The optimal solutions minimize a fidelity measure (defined in Eq. (12) for state transfer and in Eq. (14) for propagator realization) by finding a suitable time dependence of the control Hamiltonian in Eq. (20).
@@
Piece-wise control is utilized in the control Hamiltonian, dividing the pulse duration $T$ into $M$ intervals with constant driving amplitude in each interval. The optimization process aims to minimize the infidelity (Eq. (12) or Eq. (14)) by optimizing the set of control amplitudes $\{\{f_{j}^{l},h_{1}^{l},h_{n}^{l}\}:1\leq l\leq M,1\leq j\leq n\}$ and the total duration $T$. To efficiently evaluate the infidelity during optimization, the operator basis elements $\{\mathfrak{a}_{j}\}$ are normalized, i.e. satisfy $\mbox{tr}(\mathfrak{a}_{j}\mathfrak{a}_{k})=\delta_{jk}$. This normalization simplifies the propagation of operators according to Eq. (8) and the subsequent calculation of the overlap between the target coefficient vector $\bm{a}_{T}$ and the final coefficient vector $\bm{a}(T)$ obtained from an initial condition $\bm{a}(0)$.

### v2

@@ IV Realization of propagators
Preparing an eigenstate of a Hamiltonian with specific interactions has applications in computational problems [59, 60]. In quantum simulation [41], on the other hand, it is of interest to explore the dynamics induced by a given Hamiltonian. Such applications require – in addition to an initial state preparation – the realization of propagators. Propagators characterized by Hamiltonians beyond pairwise interactions, such as $H_{G}$, $H_{C}$, $H_{D}$ discussed above in Sec. III.1 cannot be realized by static means, and the present approach can be used to find time-dependent functions $f_{j}(t)$, $w_{1}(t)$ and $w_{n}(t)$ in the system Hamiltonian $H(t)$ (Eq. (20)), such that the resultant propagator is of the form $\exp(-iH_{T}T)$ with a Hamiltonian including more-than-pairwise interactions at a given time $T$.
The realization of such effective Hamiltonians is commonly pursued based on Floquet theory [61] and the present approach allows us to go beyond the common perturbative treatment that is typically required [62].
The realization of desired propagators is exemplified here in more detail for $\exp(-i\pi/8\,H_{G})$, $\exp(-i\pi/8\,H_{C})$ and $\exp(-i\pi/8\,H_{D})$ given by the three Hamiltonians $H_{G}$ (Eq. (22)), $H_{C}$ (Eq. (23)) and $H_{D}$ (Eq. (24)), and system dynamics in terms of the Hamiltonian $H(t)$ in Eq. (20).
@@ IV.1 Objective function
As shown in Sec. A of the Appendix, the only operators that commute with each of the operators in Eq. (32c) are multiples of the identity, which leaves $U_{T}^{\dagger}U=\exp(i\varphi)\mathbbm{1}$ as the only solution.
@@
Figure 4: Optimised pulses, $f_{1}$, $f_{25}$ and $f_{50}$, for the realization of propagators in a chain with 50 spins as function of time $t$ in unit of $\tau_{g}$. The resultant operator infidelities are $\approx 4\times 10^{-5}$ for $U_{C}$ (a), $\approx 9\times 10^{-5}$ for $U_{D}$ (b), and $\approx 10^{-4}$ for $U_{G}$ (c). For $U_{G}$ $f_{50}=f_{1}$ and hence is not shown. The pulses for $U_{G}$ in (f) are longer due to the difficulty in quantum-simulating an $n$-body interaction with a pairwise controlled Hamiltonian.
@@ IV.2 Results – Propagators
Fig. 2b depicts the infidelities ${\cal J}_{{\cal U}_{T}}$, defined in Eq. (19) together with Eq. (32c), as function of system size $n$. The solid red diamonds, blue squares, and black triangles correspond to the target unitaries $U_{G}=\exp(-i\pi/8\,H_{G})$, $U_{C}=\exp(-i\pi/8\,H_{C})$ and $U_{D}=\exp(-i\pi/8\,H_{D})$, respectively. This is based on optimized piecewise constant pulses with a duration of $\sim n\tau_{g}$ and $20n$ bins for $U_{G}$, and $\sim\tau_{g}$ and $10n$ bins for $U_{C}$ and $U_{D}$.
@@
Also, the duration of controlled dynamics required for the realization of a desired propagator has a fundamental limit given by the timescale $\tau_{g}=2\pi/g$. The numerically observed minimal time required to implement the unitaries induced by $H_{C}$ and $H_{D}$ is independent of the system size, but the minimal time required to implement the unitary induced by $H_{G}$ scales linearly in the number of qubits. Given that $H_{C}$ and $H_{D}$ contain at most three-body interactions, whereas $H_{G}$ contains an $n$-body interaction, i.e. an interaction whose type changes with increasing system size $n$, the numerically observed dependence of minimal durations of controlled dynamics seem consistent with the underlying problem, which generates some confidence that the obtained solutions are close to the actual global solution of the control problem.
Fig. 4 depicts optimized driving patterns $f_{1}$, $f_{25}$ and $f_{50}$ obtained with the same procedure of smoothing and reoptimization as in Fig. 3. Given the fundamental difficulty of quantum simulating the n-body interaction $\prod_{i=1}^{n}Z_{i}$ in $U_{G}$ with a system Hamiltonian including only pairwise interactions, the required duration of the controlled dynamics is substantially longer than in the cases of $U_{C}$ and $U_{D}$. Quite surprisingly, the second half of the time-period shown in panel c is essentially uncontrolled dynamics. Nevertheless, the full duration of $50\tau_{g}$ seems to be necessary in order to reach high fidelities, but the controlled dynamics during the first half seems to drive the system to a state that evolves mostly under the influence of the $XX$ interactions towards the target state in the interval $[25\tau_{g},50\tau_{g}]$. The pulse data for smaller system sizes are provided in Ref. [58].
@@ V.2 Driven spin comb
Figure 5: Spin comb with nearest-neighbor $XY$ interactions in the top chain and $ZZ$ interactions between spins of different chains (inset (a)). Single-spin $X$-driving (inset (a)) enables the realization of effective dynamics with four-spin interactions matching the Toric Code Plaquette operators Eq. (42) (inset (b)).
An interaction geometry that results in a Lie algebra of quadratic size in the qubit number $n$ that can be used to realize elementary building blocks of topologically protected quantum information processing [63] is depicted in Fig. 5(a). It is a comb-structure with a linear chain of $m=n/2$ qubits with nearest-neighbor interactions, and a second chain of another $m$ qubits interacting only with qubits on the first chain.
@@ V.3 Hexagonal spin ladder
with $m=n/2$, $j\in[1,(m-1)/2]$, $l\in[1,(m+1)/2]$. The resultant Lie algebra can be constructed systematically similarly to the discussion in Sec. V.1; because of the more intricate interaction geometry the exact form of the individual elements is a bit bulkier, and it is deferred to Appendix D.
@@
VI Discussion and Outlook
The techniques discussed in this paper integrate into and form a part of the spectrum of quantum optimal control techniques. They are inspired by control techniques that are based on analytic forms ${\cal I}_{a}(t)$ of dynamical quantum invariants. For the explicit system Hamiltonian $H(t)$ (Eq. (20)) discussed in this paper, such an analytic form ${\cal I}_{a}(t)$ would need to be such that a Hamiltonian that satisfies Eq. (1) together with ${\cal I}_{a}(t)$ contains only the terms $Z_{j}$ ($j\in[1,n]$), $X_{1}$, $X_{n}$ and the interaction term $\sum_{j=1}^{n}X_{j}X_{j+1}$, but none of the additional terms listed in Eq. (21j). Finding such an analytic expression seems to be an insurmountable task, and the main goal of this paper is to avoid the dependence on such analytic expressions.
Notable existing approaches with goals similar to those pursued in this paper include counterdiabatic driving [67, 68, 69, 70], which requires additional terms in the system Hamiltonian to ensure that system dynamics follow adiabatic evolution, even beyond the regime where the adiabatic approximation holds. For the system Hamiltonian $H(t)$ (Eq. (20)), counterdiabatic terms would typically include terms listed in Eq. (21j), but that are not contained in the system Hamiltonian in Eq. (20). Such terms can typically not be realized exactly, but they can be approximated as the result of higher-order processes in a time-dependent Hamiltonian [67]. The present approach aims at avoiding such approximations, and the existence of a closed commutation relation Eq. (5) enables doing this in a numerically exact fashion.
The realization of propagators defined by Hamiltonians that cannot be realized in practice by static means discussed in Sec. IV is commonly pursued in the context of Floquet engineering [71, 72]. While Floquet theory guarantees the existence of such propagators at periodically occurring points in time for periodically time-dependent Hamiltonians, it does not provide us with the ability to construct such propagators exactly. Existing approaches are thus typically based on perturbative constructions. Also in this context, the present approach can be helpful to overcome the restriction to perturbative regimes and to find numerically exact solutions.
The construction of explicit control solutions in this work makes use of the popular algorithm for computing the gradients with respect to control parameters utilized in GRAPE [13, 73], but the structural equivalence between the Schrödinger equation and Eq. (8) implies that the full range of optimal quantum control algorithms is applicable to the present approach. For example, more recent methods to construct gradients [74], alternative parametrizations of the control functions (e.g. in terms of Gaussian pulses [16]), or formulations that do not require back-propagation during optimization  [75] could be employed to achieve smoother solutions and potentially enhance convergence of optimal control algorithms.
While the exponential scaling of composite quantum systems poses fundamental constraints that cannot be overcome in general, the fact that there are sets of operators with favourable scaling does allow us to overcome them in specific cases. With control over interaction geometries in various platforms such as superconducting qubits [76], trapped ions and Rydberg atoms [77] the present approach can help to fully exploit the qubit count in current devices (ranging up to $256$ [77]) without the need of long gate sequences that are conflicting with decoherence and the accumulation of individual gate errors. While verification of the functionality of a quantum device is a challenging problem [78], the description of quantum dynamics in terms of invariants can also be used to make predictions for the targeted dynamics that can be experimentally tested.
@@ Acknowledgements
Appendix A Absence of symmetries in the system Hamiltonian
@@
$H(t)=\sum_{j=1}^{n}f_{j}(t)Z_{j}+g\sum_{j=1}^{n-1}X_{j}X_{j+1}+\sum_{j\in\{1,n\}}w_{j}(t)X_{j}$ (45)
for $g\neq 0$ and general time-dependent functions $f_{j}(t)$ and $w_{j}(t)$, i.e.
@@
with an $n-2$-qubit operator $D^{n}$ that can be expanded in terms of tensor-products of $\mathbbm{1}$ and $Z$ terms, or more explicitly as
@@
Appendix B Numerical simulation and optimization
@@
This section describes the numerical techniques used to find the optimal control strategies for the problems of state transfer and realization of propagators, as detailed in Secs. III and IV. The optimal solutions minimize a fidelity measure (defined in Eq. (12) for state transfer and in Eq. (19) for propagator realization) by finding a suitable time dependence of the system Hamiltonian in Eq. (20).
The operator is propagated with the equation of motion Eq. (8), where the matrix $K(t)$ is given by the structure constants (Eq. (9)). We provide the structure constants for the Lie algebra of Eq. 21j in a computer readable format in Ref. [58]. One can obtain these structure constants by computing the commutator relation between the basis elements either analytically, or numerically with the help of the Pauli Algebra package in Mathematica [82].
B.1 Coefficient vectors of the initial and target operators
In our method, numerical propagation and optimisation are done entirely with the coefficients in a basis expansion of the operators, i.e., the coefficients $a_{j}$ in the expansion
${\cal I}=\sum_{j}a_{j}\mathfrak{a}_{j}\ ,$ (55)
where $\mathfrak{a}_{j}$ are the basis elements listed in Eq. (21j). In this section, we describe how to find these coefficient vectors for the initial and target operators discussed in the paper.
B.1.1 State preparation
The initial operator $\mathcal{I}_{0}$ (Eq. (25)) and the targets $\mathcal{I}_{G}\equiv H_{G}$ (Eq. (22)) and $\mathcal{I}_{C}\equiv H_{C}$ (Eq. (23)) are already written as explicit expansions of the basis elements, so the coefficients can be readily read off.
The control target $\mathcal{I}_{D}$ with the same ground state as $H_{D}$ is obtained by evolving the initial operator ${\cal I}_{0}$ under the adiabatic Hamiltonian
$H_{a}(t)=(1-t/T_{a})\sum_{j}Z_{j}+(t/T_{a})H_{D}$
$+(t/T_{a})(1-t/T_{a})\left(\sum_{j}Z_{j}+\sum_{j=1}^{n-1}X_{j}X_{j+1}+X_{1}+X_{n}\right),$ (56)
where $T_{a}$ is a sufficiently long duration. In the $d-$dimensional space of the Lie algebra, $H_{a}(t)$ corresponds to
$K_{a}(t)=(1-t/T_{a})\sum_{j}K^{z}_{j}+(t/T_{a})K_{D}$
$+(t/T_{a})(1-t/T_{a})\left(\sum_{j}K^{z}_{j}+\sum_{j=1}^{n-1}K^{xx}_{j}+K^{x}_{1}+K^{x}_{n}\right)\ ,$ (57)
where $K^{z}_{j},K_{D},K^{xx}_{j}$, and $K^{x}_{j}$ are the $d\times d$ adjoint representation matrices of $Z_{j},H_{D},X_{j}X_{j+1}$ and $X_{j}$, respectively (see Eq. (9)). Starting with the initial coefficient vector $\bm{a}_{0}$ of $I_{0}$, one propagates with
$\dot{\bm{a}}(t)=K_{a}(t)\bm{a}(t)\ ,$ (58)
for the duration $T_{a}$ to obtain the coefficient vector $\bm{a}_{D}$ of $I_{D}$.
B.1.2 Propagators
Realizing a propagator $U_{T}$ requires more than a single pair of initial and target operators $\{\mathcal{I}(0),U_{T}\mathcal{I}(0)U_{T}^{\dagger}\}$. Instead, $2n$ pairs of initial and final operators, $\{\mathcal{I}_{j}(0),\mathcal{I}_{Tj}\equiv U_{T}\mathcal{I}_{j}(0)U_{T}^{\dagger}\}$, are sufficient, with the initial operators $\mathcal{I}_{j}(0)$ selected from Eq. (32c). Obtaining the coefficient vectors of $\mathcal{I}_{j}(0)$ is straightforward as they are already expressed in terms of the basis elements of the Lie algebra in Eq. (21j).
The propagators $U_{G}$, $U_{C}$ and $U_{D}$ considered in our work are of the common form
$U_{T}=e^{-i\theta\mathcal{H}}\ ,$ (59)
where $\theta$ is a phase, and $\mathcal{H}$ a linear combination of the basis elements of the Lie algebra. The final condition $\mathcal{I}_{Tj}\equiv U_{T}\mathcal{I}_{j}(0)U_{T}^{\dagger}$ corresponds to propagating $\mathcal{I}_{j}(0)$ with the Hamiltonian $H(t)$ (Eq. (20)) for a duration $\theta$. Starting with the coefficient vector $\bm{a}(0)$ of $\mathcal{I}_{j}(0)$, one propagates the equation
$\dot{\bm{a}}(t)=K\bm{a}(t)\ ,$ (60)
where $K$ is the adjoint representation matrix of $H(t)$, for the duration $T_{P}=\theta$. Then, $\bm{a}(T_{P})$ is the coefficient vector of the target $\mathcal{I}_{Tj}$.
B.2 Computing infidelity and gradient
@@
Piece-wise control is used for the individual $X_{j}$ and $Z_{j}$ drivings of the system Hamiltonian (Eq. (20)), while the coupling strength $g$ in the interaction term $X_{j}X_{j+1}$ is assumed to be constant. The pulse duration $T$ is divided into $M$ intervals, each with a fixed driving amplitude. The optimization process aims to minimize the infidelity (Eq. (12) or Eq. (19)) by optimizing the set of control amplitudes $\{\{f_{j}^{l},w_{1}^{l},w_{n}^{l}\}:1\leq l\leq M,1\leq j\leq n\}$ and the total duration $T$. To efficiently evaluate the infidelity during optimization, the operator basis elements $\{\mathfrak{a}_{j}\}$ are normalized, i.e. satisfy $\mbox{tr}(\mathfrak{a}_{j}\mathfrak{a}_{k})=\delta_{jk}$. This normalization simplifies the propagation of operators according to Eq. (8) and the subsequent calculation of the overlap between the target coefficient vector $\bm{a}_{T}$ and the final coefficient vector $\bm{a}(T)$ obtained from an initial condition $\bm{a}(0)$.
The control problem is formulated within the Lie algebra as finding the final coefficient vector $\bm{a}(T)$ to minimize the invariant infidelity $\mathcal{J}=1-\bm{a}_{T}\cdot\bm{a}(T)$, where $\bm{a}_{T}$ is the target coefficient vector, starting from an initial condition $\bm{a}(0)$. The evolution of $\bm{a}(t)$ is governed by the adjoint representation matrix $K(t)$ of the system Hamiltonian $H(t)$. If one expresses $H(t)=H_{0}+\sum_{k}c_{k}(t)H^{c}_{k}$, where $c_{k}(t)$ is the amplitude of the control operator $H^{c}_{k}$, then $K(t)=K_{0}+\sum_{k}c_{k}(t)K^{c}_{k}$ where $K_{0}$ and $K^{c}_{k}$ are the adjoint representations of $H_{0}$ and $H^{c}_{k}$, respectively (see Eq. (9)). For the piece-wise control defined above, the solution of $\dot{\bm{a}}(t)=K(t)\bm{a}(t)$ is
$\bm{a}(T)=\prod_{l=M}^{1}\mathcal{U}_{l}\bm{a}(0)\ ,$ (61)
where
$\mathcal{U}_{l}=e^{\Delta t\left(K_{0}+\sum_{k}c_{k}^{l}K_{k}^{c}\right)}\ .$ (62)
Note that $\mathcal{U}_{l}$ is unitary as $K(t)$ is anti-hermitian. One first computes and stores the set of forward and backward propagated vectors, defined by
$\bm{a}_{m}$ $=\mathcal{U}_{m}\mathcal{U}_{m-1}\dots\mathcal{U}_{1}\bm{a}(0)\ ,$
$\bm{b}^{\dagger}_{m}$ $=\bm{a}_{T}^{\dagger}\mathcal{U}_{M}\mathcal{U}_{M-1}\dots\mathcal{U}_{m}\ ,$
using the recursive relations $\bm{a}_{m}=\mathcal{U}_{m}\bm{a}_{m-1}$ and $\bm{b}^{\dagger}_{m}=\bm{b}^{\dagger}_{m+1}\mathcal{U}_{m}$, then the gradient of $\mathcal{J}$ with respect to $c^{l}_{k}$ is
$\frac{\partial\mathcal{J}}{\partial c_{k}^{l}}=\frac{\partial}{\partial c_{k}^{l}}\left(1-\bm{b}^{\dagger}_{l+1}U_{l}\bm{a}_{l-1}\right)=-\bm{b}^{\dagger}_{l+1}\frac{\partial U_{l}}{\partial c_{k}^{l}}\bm{a}_{l-1}\ .$ (63)
The derivative of $\mathcal{U}_{l}$ is approximated by a 2nd-order method [83]
$\frac{\partial U_{l}}{\partial c_{k}^{l}}=\left\{\Delta tK^{c}_{k}-\frac{\Delta t^{2}}{2}\left[K_{l},K^{c}_{k}\right]\right\}U_{l}+O(\Delta t^{3})\ ,$ (64)
where $K_{l}=K_{0}+\sum_{k}c_{k}^{l}K^{c}_{k}$, thus
$\frac{\partial\mathcal{J}}{\partial c_{k}^{l}}=-\bm{b}^{\dagger}_{l+1}\left\{\Delta tK^{c}_{k}-\frac{\Delta t^{2}}{2}\left[K_{l},K^{c}_{k}\right]\right\}\bm{a}_{l}\ ,$ (65)
which is computed by matrix-vector multiplication. The most computationally expensive part is the multiplication of a matrix exponential with a vector for obtaining $\bm{a}_{m}$ and $\bm{b}^{\dagger}_{m}$, but it can be done efficiently with the Krylov subspace method owing to the sparsity of $K(t)$.
For calculating the derivative of $\mathcal{J}$ with respect to the duration $T$, one simply calculates $\mathcal{J}$ again for a perturbed duration $T+dT$ while keeping the amplitudes $c_{k}^{l}$ the same, so that
$\frac{\partial\mathcal{J}}{\partial T}\approx\frac{\mathcal{J}(T+dT)-\mathcal{J}(T)}{dT}\ .$ (66)
In our calculation, we choose $g\,dT=10^{-10}$.
B.3 Optimisation